INAF queue observing, 2016 July 2 - 3

Observer: F. Cusano, A. Rossi
Telescope Operator: G. Bechetti
Support Astronomer: O Kuhn on call
Instrument: MODS, LUCI

Summary

At the sunset the sky is clear.
Problem with the Dx AdSec, we can not use MODS in binocol mode and we can not use LUCI2 also.
We use MODS1.

Night duration (12\xB0 el twilight) 20:30-4:00 ( 03:30-11.00 UT) = 7h 30m

Observing time: 7.2 (95%)

Weather time loss: 0 h (00%)

technical time loss 0.3h (5%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
Della Ceca 2XMMJ135055.7+642857 dual grating 2h completed
Testa J1935+1726 dual grating 3h completed
Calib slitflat_5as dualgrating    
Standards   dual grating    

Details

2:30 UT the TO opens the dome

3:50 waiting for AdDec to ricover

4:04 still problem with AdDec, we decide to use just MODS1

4:22 slew to Della Ceca, seeing 1", acqMODS Target_11B.acq mods1r.20160703.0002-0004

4:38 execMODs mods1r.20160703.0005-0008, mods1b.20160703.0002-0005
  • seeing 1" at 4:40, humidity 33%
  • seeing 0.8" at 5:28

6:48 slew to standard BD+33 2642, dual grating
  • acqMOD mods1r.20160703.0009-0010

6:53 execMODS mods1r.0011-0012,mods1b.006-007
  • 7:05 the telescope is not tracking we stop the integration, spectra 2 out of 3 good

07:09 we slew to Testa_A, acqMODS 2015B_3_A.acq mods1r.0014-0016
  • seeing 1"
  • 7:15 changing the guide star it was a double

07:38 execMODs 2015B_3_A_1.obs, mods1r.0017-19,mods1b.0008-10
  • seeing 1.1" at 7:40

08:46 execMODs 2015B_3_A_2.obs, mods1r.0020-22,mods1b.0011-13
  • seeing 1.0" at 9:00

09:54 execMODs 2015B_3_A_3.obs, mods1r.0023-0025,mods1b.0014-0016
  • seeing 1.0" at 10:00
  • passing cirrus since 10:15 UT
  • seeing 1.3" at 10:20
  • seeing 1.1" at 10:40

11:01 acqMODS BD284211 mods1r.0026-27

11:09 execMODs BD284211_LSx60x5.obs, mods1r.0028-0030,mods1b.0017-0019

11:25 END OF NIGHT DOME CLOSED

11:30 MODS Calibrations slitflat for standard
  • execMODS slitflat_5as.cal mods1r.0031-,mods1b.0020-

11:30 LBC bias

-- %USERSIG{AndreaRossi - 2016-07-03}%

Comments

 
Topic revision: r12 - 15 Dec 2017, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback